I would appreciate it if you can help me this is a Quadratic function question.
I would try to begin but I have no idea on how to start.


1) A ball is thrown upward with an initial velocity of 32 ft./sec from a height of six ft.
The function s(t)=-16t*+32t+6 gives the height of the ball t seconds after release. Determine the time when the ball is at maximum height, and find that height.

2) Jessica can be paid one of two ways for painting a house:
Plan A: $200 plus $12 per hour
Plan B: $20 per hour
Suppose the job takes n hours to complete. For what values of n is Plan A better for Jessica?

3) In a study, it was found that the threshold weight, W, defined as the weight above which the risk of death rises dramatically, is given by
W(h)=( h/12.3) cube
Find the threshold weight for someone 5 ft 7 inches tall.

Thank you for your help!

Answer by feliz1965(151)

It is better to place one question at a time if you want tutors to answer your questions here.
I will do number 3.
3) In a study, it was found that the threshold weight, W(h), defined as the weight above which the risk of death rises dramatically, is given by
W(h)=( h/12.3)^3
Find the threshold weight for someone 5 ft 7 inches tall.
There is a typo is your question. The threshold weight cannot be simply W.
The weight we need to find is defined as a fucntion and so what we need to find W(h), which is read "W of h."
To do so, replace h with 5 feet and 7 inches and simplify the cubic function using a calculator.
I will convert feet to inches for easy reading.
So, 5 feet = 60 inches.
Then 60 + 7 = 67 inches.
W(67 inches) = (67 inches/12.3)^3
W(67 inches) = (5.45)^3
W(67 inches) = 161.9
So, the threshold weight, according to your numbers, is 161.9